OpenStudy (swaugh):

medal + fan awaiting What is the length of line segment A B if A is point one comma one and B is point nine comma seven? The point A is plotted at one comma one and point B is plotted at nine comma seven on a coordinate plane. The points are connected with a diagonal line segment. 8 ten twelve fourteen

OpenStudy (zpupster):

use the distance formula\[\sqrt{(x2-x1)^2+(y2-y1)^2 }\] (x1,y1) =(1,1) (x2,y2)=(9,7)

OpenStudy (zpupster):

because i am practiced with geogebra